OpenStudy (anonymous):

Find the amount A accumulated after investing a principal P=$17,000 for t=9 years at an interest rate r=4% compounded monthly. (Assume no withdrawals are made.) The amount after 9 years will be $____

OpenStudy (anonymous):

FV = PV(1+i/12)^n(12) FV = 17000(1+ 0.04/12)^9(12) FV = 24352.02

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

np glad to help